OpenStudy (anonymous):

help write an quadratic equation having the given numbers as solutions -10 and -1

OpenStudy (anonymous):

if a,b are the solutions to a quadratic then the quadratic is given by (x-a)(x-b)=0

OpenStudy (anonymous):

thus here it must be the product of (x+1)(x+10)=0

OpenStudy (anonymous):

This is saying if x=-10 then \[ax^2+bx+c= 0\] and the same thing for -1 we know we can factor this into \[(x \pm d )(x \pm e)\] so if either of those are zero then the whole thing is zero and we call those numbers roots. They gave of us roots so we get \[(x+10)(x+1)=x^2+11x+10\] you can check this by plugging in the numbers given

OpenStudy (raden):

alternative : x^2 - (a+b)x + ab = 0 (with a, b are its solution) here given a=-10 and b=-1 (or reverse) x^2 - (-10-1)x + (-10)(-1) = 0 x^2 + 11x + 10 = 0